Circular No. 6910

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 980519 J. M. Muller, BeppoSAX Science Data Center, Rome, and Space Re-

search Organization of the Netherlands (SRON), Utrecht; J. Heise, SRON; C. Butler, Agenzia Spaziale Italiana, Rome; F. Frontera, Istituto Tecnologie e Studio Radiazioni Extraterrestri (ITESRE), Bologna; L. Di Ciolo, G. Gandolfi and A. Coletta, BeppoSAX Science Operation Center, Telespazio, Rome; and P. Soffitta, Istituto Astrofisica Spaziale (IAS), CNR, Frascati, communicate on behalf of the BeppoSAX team: “On May 19.51410 UT, a γ-ray burst (Bacodine trigger 6764) was detected by the Gamma-Ray Burst Monitor onboard BeppoSAX. A preliminary analysis shows a complex burst lasting 30 s with a peak intensity of ∼ 1500 counts/s (40–700 keV). An x-ray counterpart of this burst was detected by the Wide Field Camera WFC2 having a duration of 190 s and a peak flux of ∼ 2.5 Crab (2–28 keV). The position derived from the WFC2 image is α = 23^{h}22^{m}15^{s}, δ = +77^{o}15′.0 (equinox 2000.0; error radius 3′). A BeppoSAX follow-up observation is in progress.”

GRB 980515 R. Ricci and G. D’Andreta, BeppoSAX Science Operation Center,

Rome; E. Costa and P. Soffitta, IAS; J. in ’t Zand, SRON; J. M. Muller, BeppoSAX Science Data Center and SRON; L. Nicastro, Istituto di Fisica Cosmica ed Applicazioni dell’Informatica, CNR, Palermo; and E. Palazzi, ITESRE, report: “The position of GRB 980515, observed by the BeppoSAX Gamma-Ray Burst Monitor and Wide Field Camera (WFC, unit 2) on May 15.708495 UT, has been refined by off-line data analysis. The final position is α = 21^{h}17^{m}25^{s}, δ = −67^{o}15′.3 (equinox 2000.0), with an error radius of 4′. The event duration was ∼ 15 s in γ-rays (in the energy band 40–700 keV), with a peak intensity of 450 counts/s. The peak intensity in the WFC (in the energy band 2–28 keV) was ∼ 1.5 Crab. The source 1SAX J2116.8–6712 (IAUC 6909) is therefore positionally consistent with GRB 980515.”

COMET C/1998 J1 (SOHO) R. H. McNaught, Bugaldie, N.S.W., reports that a photograph taken

on May 19.36 UT with a 55-mm f/1.4 camera lens and HP5 film reveals a gas tail 10^{o} long in p.a. 60^{o} and a dust tail just over 1^{o} long in p.a. 80^{o}. At the same time, McNaught viewed the comet via naked eye (m_{1} ∼ 4).

1998 May 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green